Let f(x)=x^3-6x^2+3x+10, then choose the set of correct options regarding f(x).


  • If x element of (-\infinity, 1] union (2, 5) x∈(−∞,1]∪(2,5), then f(x) is negative.


  • If x element of [-2, 2] union (5, \infinity)x∈[−2,2]∪(5,∞), then f(x) is positive.


  • If x element of [0, 1] union (10, \infinity)x∈[0,1]∪(10,∞), then f(x) is positive.


  • If x element of [2, 3] union (5, \infinity) x∈[2,3]∪(5,∞), then f(x) is positive.


  • If x element of (-1, 2) union (5, \infinity) x∈(−1,2)∪(5,∞), then f(x) is positive.
1
Expert's answer
2021-10-05T17:55:18-0400
"f(x)=0=>x^3-6x^2+3x+10=0"

"x^2(x+1)-7x(x+1)+10(x+1)=0"

"(x+1)(x^2-7x+10)=0"

"(x+1)(x-2)(x-5)=0"

"x_1=-1, x_2=2, x_3=5"

If "x<-1," then "f(x)<0."

If "-1<x<2," then "f(x)>0."

If "2<x<5," then "f(x)<0."

If "x>5," then "f(x)>0."


Answer:

5. If "x\in(-1, 2)\cup(5,\infin)," then "f(x)" is positive.